1. What mass of zinc is needed to react with 23.1 g of phosphoric acid according to the following equation?

3 Zn    + 2 H3(PO4)    🡪  3  H2    +    1 Zn3(PO4)2




1
Expert's answer
2022-05-16T16:51:04-0400

According to the equation, n (Zn) = 3/2 x n (H3PO4)

n = m / M

M (H3PO4) = 98 g/mol

M (Zn) = 65.4 g/mol

n (H3PO4) = m / M = 23.1 / 98 = 0.24 mol

n (Zn) = 3/2 x n (H3PO4) = 3/2 x 0.24 = 0.36 mol

m (Zn) = n x M = 0.36 x 65.4 = 23.5 g